# On the Dirichlet problem for the degenerate $k$-Hessian equation

### v1

@@ 4. Reduction of the second-order estimate to the boundary
where $K$ depends on $\Omega$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$. The proof of Theorem 1.2 is complete. ∎
@@ 5. The weakly interior estimate
$\hskip 12.80373pt+\frac{1}{(\psi+\beta)^{\alpha+1}}\left(\frac{\alpha-1}{4}|\xi|^{2}\psi_{x_{i}}\psi_{x_{j}}-2\alpha\psi_{\xi}\psi_{\xi x_{i}}\psi_{x_{j}}-2\alpha\psi_{\xi}\psi_{x_{i}}\psi_{\xi x_{j}}-\alpha\psi_{\xi}^{2}\psi_{x_{i}x_{j}}\right)$
@@
$\hskip 12.80373pt+\frac{G^{ij}}{(\psi+\beta)^{\alpha+1}}\left(\frac{\alpha-1}{4}|\xi|^{2}\psi_{x_{i}}\psi_{x_{j}}-(1+\alpha)\psi_{\xi}\psi_{x_{i}}\xi_{j}+3\alpha\psi_{\xi}^{2}\psi_{x_{i}x_{j}}+\frac{1}{2}\alpha\psi_{\xi}^{2}\delta_{ij}\right)$
@@
$\leq\frac{1}{(\psi+\beta)^{\alpha}}\left(-\frac{1-\alpha}{4\alpha}+C_{1}\|\psi\|_{C^{3}(\overline{\Omega})}^{2}\right)|\xi|^{2}$
$\hskip 12.80373pt+\frac{G^{ij}}{(\psi+\beta)^{\alpha}}\left[-\frac{1}{4\alpha}\xi_{i}\xi_{j}-\frac{1+\alpha}{\psi+\beta}\psi_{\xi}\psi_{x_{i}}\xi_{j}-\frac{\alpha^{2}+3\alpha}{(\psi+\beta)^{2}}\psi_{\xi}^{2}\psi_{x_{i}}\psi_{x_{j}}\right],$ (5.10)
@@
$=\sum_{i,j=1}^{n}G^{ij}\left(\sqrt{\alpha^{2}+3\alpha}\frac{\psi_{\xi}}{\psi+\beta}\psi_{x_{i}}+\frac{1+\alpha}{2\sqrt{\alpha^{2}+3\alpha}}\xi_{i}\right)\left(\sqrt{\alpha^{2}+3\alpha}\frac{\psi_{\xi}}{\psi+\beta}\psi_{x_{j}}+\frac{1+\alpha}{2\sqrt{\alpha^{2}+3\alpha}}\xi_{j}\right)$
@@
$\mathscr{L}\big((\psi+\beta)\eta^{2}\big)$ $=G^{ij}\left(\partial_{x_{i}x_{j}}+2q_{j}\partial_{\eta x_{i}}+q_{i}q_{j}\partial_{\eta\eta}\right)\left[(\psi+\beta)\eta^{2}\right]$
@@
$\mathscr{L}(w)$ $=G^{ij}\left(\partial_{x_{i}x_{j}}+2r\partial_{x_{i}\xi_{j}}+2q_{j}\partial_{x_{i}\eta}+r^{2}\partial_{\xi_{i}\xi_{j}}+2rq_{j}\partial_{\xi_{i}\eta}+q_{i}q_{j}\partial_{\eta\eta}-2q_{j}\partial_{\xi_{i}}\right)[w]$
@@
$\geq-K_{1}\sigma_{1}^{\frac{-1}{k-1}}\left(6|\xi|^{2}+\eta^{2}+r^{2}\right)-K_{1}\left(4\eta^{2}+r^{2}\right)-4r^{2}-\left(\|u\|_{C^{1}(\overline{\Omega})}^{2}+2\|u\|_{C^{0}(\overline{\Omega})}\right)|q|^{2}$
$\geq-K_{1}\sigma_{1}^{\frac{-1}{k-1}}\left(6|\xi|^{2}+\eta^{2}+\frac{C_{\psi}}{\beta^{2}}|\xi|^{2}\right)-K_{1}\left(4\eta^{2}+\frac{C_{\psi}}{\beta^{2}}|\xi|^{2}\right)-4\frac{C_{\psi}}{\beta^{2}}|\xi|^{2}$
$\hskip 12.80373pt-\left(\|u\|_{C^{1}(\overline{\Omega})}^{2}+2\|u\|_{C^{0}(\overline{\Omega})}\right)\frac{C_{\psi}}{\beta^{4}}|\xi|^{2}$
@@
$\sigma_{1}\big[\lambda\big(D^{2}u\big)\big]\geq\lambda_{\operatorname{max}}\big(D^{2}u\big)\geq u_{\xi_{*}\xi_{*}}\geq\frac{1}{\beta^{3}}-2\eta_{*}u_{\xi_{*}}-\eta_{*}^{2}u\geq\frac{1}{\beta^{3}}-3\|u\|_{C^{1}(\overline{\Omega})}>1,$ (5.19)
@@
$=\left(\frac{1}{4\alpha}(\psi+\beta)^{1-\alpha}+\frac{\psi_{\xi}^{2}}{(\psi+\beta)^{\alpha}}\right)\left(\beta^{\alpha}\sup_{\partial\Omega}u_{\nu\nu}+\frac{4K+2C}{\beta^{8}}\right)$ (5.20)
@@
using Theorem 1.1, we complete the proof of Theorem 1.5. ∎

### v2

@@ 3. Reduction to the boundary second-order normal estimate
where $K$ depends on $\Omega$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$. The proof of Lemma 3.3 is complete. ∎
Combining (3.1), Lemmas 3.1, 3.2 and 3.3, we reduce the $C^{2}$ estimate to the boundary second-order normal derivative estimate under the condition $f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}})$ with $f\geq 0$ in $\Omega_{0}$. Under the homogeneous boundary condition, Dong [6] established the required boundary second-order normal derivative estimate. Thus the $C^{2}$ estimate independent of $\inf_{\Omega_{0}}f$ is obtained. Then the existence is established by an approximation argument.
Theorem 3.4.
We emphasize here that results corresponding to Lemmas 3.1, 3.2, 3.3, and Theorem 3.4 were established by Jiao-Wang [17] under the condition $f^{1/(k-1)}\in C^{1,1}(\overline{\Omega})$ with $f\geq 0$ in $\Omega$.
4. Monge-Ampère equation
In this section, we assume
$f^{\frac{3}{2n-2}}\in C^{2,1}(\overline{\Omega}),\quad f\geq 0\ \text{in}\ \Omega,$
and apply the method of Guan-Trudinger-Wang [12] to establish the boundary second-order normal derivative estimate. We emphasize here that for the Monge-Ampère equation, Lemmas 3.1 and 3.2 hold under the weaker condition $\inf_{\Omega}f>0$ and $f^{3/(2n-2)}\in C^{1,1}(\overline{\Omega})$. Indeed, Lemma 3.1 follows directly from the convexity of the admissible solution, and Lemma 3.2 can be adapted by differentiating with respect to a tangential vector filed, with the same argument as in the proof of later Lemma 4.3.
$0<u_{11}(0)\leq u_{22}(0)\leq\dots\leq u_{n-1,n-1}(0).$
By subtracting an affine function, let
$u(0)=0\quad\text{and}\quad\nabla u(0)=0,$
which yields by the convexity of $u$,
$u(x)\geq 0,\quad\forall x\in\overline{\Omega}.$
For convenience, denote
$b_{i}:=u_{ii}(0),\quad\forall 1\leq i\leq n-1.$
By Taylor expansion with respect to $x^{\prime}$ at the origin, we have
on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$.
Lemma 4.1 ([12]).
$h(t)=\alpha t^{2}+\beta t^{3}+R(t),\quad\forall t\in\mathbb{R},$
$|\beta|\leq(1+\gamma)\sqrt{\alpha}.$
Proof.
4.1. Lower bounds for boundary second-order tangential derivatives
Lemma 4.2.
There exists a constant $\gamma_{0}>0$ such that
$\prod_{i=1}^{n-1}b_{i}\geq\gamma_{0}f(0),$
where $\gamma_{0}$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $\|f^{3/(2n-2)}\|_{C^{2,1}(\overline{\Omega})}$, but is independent of $\inf_{\Omega}f$.
Proof.
Step 1. Define
$M:=\frac{1}{b_{1}}.$
Make a dilation $x\rightarrow y=T(x)$ defined by
$y_{i}=M_{i}x_{i}\quad\text{for}\ i=1,2,\dots,n,$ (4.3)
where
Without loss of generality, we assume $M\gg 1$, which depends on the $C^{3,1}$-norm of $\partial\Omega$. Thus, the transformation $T$ can be regarded as a blow-up transformation near the origin. Define
$v(y):=M^{2}u(x).$
Then
Step 2. Define a set
$\omega:=\left\{y\in\widetilde{\Omega}:\ y_{n}<1,\ |y_{i}|<1,\ \,\forall 2\leq i\leq n-1\right\}.$
and
[…]
where $C_{2}$ depends only on $\|\varphi\|_{C^{3,1}(\partial\Omega)}$.
Step 3. We claim that there exists a constant $K>0$, independent of $M$, such that for any ball $B_{\frac{r_{\star}}{4n}}(y_{*})$ contained in $B_{r_{\star}}(y_{\star})$,
$\inf_{B_{\frac{r_{\star}}{4n}}(y_{*})}\tilde{f}<K.$ (4.9)
Suppose, for contradiction, that (4.9) does not hold. Then
$\det D^{2}v\geq K\quad\text{in}\ B_{\frac{r_{\star}}{4n}}(y_{*}).$
Since (4.8) the function
$\frac{K^{1/n}}{2}\left((y-y_{*})^{2}-\frac{r_{\star}}{4n}\right)+C_{2}$
is an upper barrier of $v$ in $B_{\frac{r_{\star}}{4n}}(y_{*})$. Thus
$v(y_{*})\leq-K^{\frac{1}{n}}\frac{r_{\star}}{8n}+C_{2}<0$
Define
$\tilde{g}:=\tilde{f}^{\frac{3}{2n-2}}\quad\text{in}\ \widetilde{\Omega}$.
We have
$\left\|D_{y}^{3}\tilde{g}\right\|_{L^{\infty}(\widetilde{\Omega})}$ $\leq\left(\prod_{i=1}^{n-1}b_{i}\right)^{\frac{-3}{2n-2}}b_{1}^{\frac{3}{2}}\left\|D_{x}^{3}\left(f^{\frac{3}{2n-2}}\right)\right\|_{L^{\infty}(\Omega)}$
$\leq\left\|f^{\frac{3}{2n-2}}\right\|_{C^{2,1}(\overline{\Omega})}\leq C.$ (4.10)
From (4.9), there is a point $\tilde{y}\in B_{\frac{r_{\star}}{4n}}(y_{\star})$ satisfying $\tilde{f}(\tilde{y})\leq K$. The fact
yields
Combining (4.1) and (4.11) yields that there exists a constant $K_{1}$ such that
that is $\tilde{g}$ is semi-convex in $\omega$. Define
$\bar{g}:=\tilde{g}+\frac{K_{1}}{2}|y|^{2}\quad\text{in}\ \overline{\omega}.$
$\inf_{B_{\frac{r_{\star}}{4n}}(y_{*})}\bar{g}<\overline{K}$ (4.12)
for any ball $B_{\frac{r_{\star}}{4n}}(y_{*})$ contained in $B_{r_{\star}}(y_{\star})$. By (4.12) and the convexity of $\bar{g}$, it follows (see Figure 1) that
and then by the mean value theorem,
$\|\tilde{g}\|_{C^{2,1}(\overline{\omega})}\leq K_{2},$
Therefore,
$K_{2}^{\frac{2(n-1)}{3}}\geq\tilde{f}(0)=\frac{f(0)}{\prod_{i=1}^{n-1}b_{i}}.$
This completes the proof of Lemma 4.2. ∎
4.2. Lower bounds for boundary tangential-normal derivatives
Lemma 4.3.
$\left|u_{\nu\eta}(x)\right|\leq C\sqrt{b_{n-1}},$
where $\eta$ is any unit tangential vector to $\partial\Omega$ at $x$, and $C>0$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $\|f^{3/(2n-2)}\|_{C^{2,1}(\overline{\Omega})}$, but is independent of $\inf_{\Omega}f$.
Proof.
Define
$M:=\frac{1}{b_{n-1}}.$
Make a dilation $x\rightarrow y=T(x)$ defined by
$y^{\prime}=\sqrt{M}x^{\prime}\quad\text{and}\quad y_{n}=Mx_{n}.$ (4.13)
Without loss of generality, we assume $M\gg 1$, which depends on the $C^{3,1}$-norm of $\partial\Omega$. Thus, the transformation $T$ can be regarded as a blow-up transformation near the origin. Define
$v(y):=M^{2}u(x).$
Then
Define a set
$\omega:=\left\{y\in\widetilde{\Omega}:\ y_{n}<1\right\}.$
Looking at $\partial\omega\cap\{y_{n}=1\}$, by (4.14) we have
where $C_{2}$ depends only on $\|\varphi\|_{C^{3,1}(\partial\Omega)}$.
Define
$\tilde{g}:=\tilde{f}^{\frac{3}{2n-2}}\quad\text{in}\ \widetilde{\Omega}$.
By the same argument as in Step 3 of the proof of Lemma 4.2, we get
$\left\|\tilde{g}\right\|_{C^{2,1}(\overline{\omega})}\leq K,$
in $\overline{\omega}$. Consider functions
$|v_{nl}(0)|\leq C,\quad\forall 1\leq l\leq n-1,$
and similarly
We proceed from Lemma 4.3 by induction to obtain a refinement. Define
Assumption 4.4.
where constants $\theta_{i}$ and $C$ depend on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $|f^{3/(2n-2)}|_{C^{2,1}(\overline{\Omega})}$, but is independent of $\inf_{\Omega}f$.
Lemma 4.5.
where constants $\theta_{k}$ and $C$ depend on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $|f^{3/(2n-2)}|_{C^{2,1}(\overline{\Omega})}$, but are independent of $\inf_{\Omega}f$.
Proof.
Define
$M:=\frac{1}{b_{k}}.$
Make a dilation $x\rightarrow y=T(x)$ defined by
$y_{i}:=M_{i}x_{i}\quad\text{for}\ i=1,2,\dots,n,$ (4.19)
where
Without loss of generality, we assume $M\gg 1$, which depends on the $C^{3,1}$-norm of $\partial\Omega$. Thus, the transformation $T$ can be regarded as a blow-up transformation near the origin. Define
$v(y):=M^{2}u(x).$
Then
Denote
$d_{i}=\frac{b_{k}}{b_{i}}\ (\leq 1)\ \ \forall k+1\leq i\leq n-1,\ \ \big|D^{3}\tilde{\rho}\big|\leq\frac{C_{1}}{\sqrt{M}},\ \ \big|D^{4}\tilde{\rho}\big|\leq\frac{C_{1}}{M},$
Define a set
$\omega:=\left\{y\in\widetilde{\Omega}:\ y_{n}<\beta^{2},\ |y_{i}|<\beta,\ \,\forall k+1\leq i\leq n-1\right\},$
and
[…]
Using the convexity of $v$ and (4.21), we derive
By the same argument as in Step 3 of the proof of Lemma 4.2, we get
$\left\|\tilde{g}\right\|_{C^{2,1}(\overline{\omega})}\leq C_{3},$
Step 1: we prove $|Dv|\leq C$ in $\omega$. Define
We now verify that
$\underline{v}\leq v\quad\text{on}\ \partial\omega.$ (4.24)
by taking $\beta$ small. By (4.20), we have
on $\partial_{1}\omega$, since $M\gg 1$. Hence
$\underline{v}\leq\frac{1}{2}\sigma\big|y^{\prime}\big|^{2}-\frac{1}{2}Ky_{n}\leq\left(\frac{1}{2}\sigma+\frac{1}{2}K\beta\right)\left|\tilde{y}\right|^{2}+\left(\frac{1}{2}\sigma-\frac{1}{8}K\right)\left|\hat{y}\right|^{2}\leq v$
on $\partial_{1}\omega$, by taking $\sigma<1/4$ and $\beta$ small depending on $K$. It holds that
by taking $\sigma$ small depending on $\varepsilon_{0}$.
We claim that
It follows from (4.21) that
[…]
and
From (4.20), we deduce
Then $\delta=O(\beta^{3})$. Since $\widetilde{\Omega}$ is convex,
$\frac{1}{2}|p^{*}-p|\leq|\bar{p}-p^{*}|.$ (4.26)
to deduce $t=O(\beta)$. Thus
Since $|\nabla u(x)-\nabla u(0)|=O(|x^{\prime}|)$ and $|u_{\nu}(x)-u_{n}(x)|=O(|x^{\prime}|^{2})$ on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$ (see later (6.7), (6.8)), using Assumption 4.4, we infer
$\left|v_{y_{n}}(y)\right|$ $=\frac{1}{b_{k}}\left(\left|u_{\nu}\big(x^{\prime},\rho\big(x^{\prime}\big)\big)\right|+O\Big(\big|x^{\prime}\big|^{2}\Big)\right)$
$\leq C_{4}$
$v(p^{*})$ $\geq v\big(p^{0}\big)+\nabla v\big(p^{0}\big)\cdot\big(p^{*}-p^{0}\big)$
$\geq\widetilde{\varphi}\big(p^{0}\big)-C_{4}\bar{p}_{n}$
$\geq\frac{1}{2}\beta^{2}-2C_{4}\varepsilon_{0}\beta^{2}+O\big(\beta^{3}\big)$
$\geq\frac{1}{3}\beta^{2}$ (4.29)
$v(\bar{p})=\widetilde{\varphi}(\bar{p})$ $=\frac{1}{2}\sum_{i=1}^{k}\tilde{b}_{i}\left|\bar{p}_{i}\right|^{2}+\widetilde{\varphi}\big(p^{0}\big)+O\big(\beta^{3}\big)$
$\leq v(p^{*})+C_{4}\bar{p}_{n}+4\varepsilon_{0}\beta^{2}$
$\leq v(p^{*})+(2C_{4}+4)\varepsilon_{0}\beta^{2}$
$\leq\frac{9}{8}v(p^{*})$ (4.30)
In conclusion, claim (4.25) is valid.
We deduce from (4.25) that
$\underline{v}\leq v\quad\text{in}\ \overline{\omega}.$
Since $\underline{v}(0)=v(0)=0$,
$v_{n}(0)\geq\underline{v}_{n}(0).$
Using the convexity of $v$ and (4.22), we get
$|Dv|\leq C\quad\text{in}\ \overline{\omega}$
Step 2. Define tangential vector fields
in $\overline{\omega}$. Indeed, for any $l\in\{1,2,\dots,n-1\}$,
Construct functions
From (4.33), at the point $y_{*}$,
$0$ $\geq F^{ij}_{z}\big[D^{2}_{z}v\big]\Psi^{l}_{z_{i}z_{j}}$
$\geq\pm F^{ij}_{z}\big[D^{2}_{z}v\big]\left(\tilde{\eta}^{l}_{z}\cdot\nabla_{z}(v-\widetilde{\varphi})\right)_{z_{i}z_{j}}+2A\operatorname{tr}F^{ij}\big[D^{2}v\big]$
$\hskip 12.80373pt+B_{A}\bigg(\frac{1}{2}F\big[D^{2}\underline{v}\big]-F\big[D^{2}v\big]\bigg)$
$\geq\pm F^{ij}_{z}\big[D^{2}_{z}v\big]\left(\tilde{\eta}^{l}_{z}\cdot\nabla_{z}v_{z_{i}z_{j}}+2\partial_{z_{j}}\tilde{\eta}^{l}_{z}\cdot\nabla_{z}v_{z_{i}}\right)+(2A-C)\operatorname{tr}F^{ij}$
$=\pm\tilde{\eta}^{l}\cdot\nabla\left(\tilde{g}^{\frac{p}{n}}\right)\pm 2\sum_{i=1}^{n}F^{ii}_{z}\big[D^{2}_{z}v\big]v_{z_{i}z_{i}}\times\partial_{z_{i}}\tilde{\eta}^{l,i}_{z}+(2A-C)\operatorname{tr}F^{ij}$
$\geq-\left|\frac{p}{n}\tilde{g}^{\frac{p}{n}-\frac{1}{2}}\frac{\tilde{\eta}^{l}\cdot\nabla\tilde{g}}{\sqrt{\tilde{g}}}\right|-C\sum_{i=1}^{n}\left|F^{ii}_{z}\big[D^{2}_{z}v\big]v_{z_{i}z_{i}}\right|+(2A-C)\operatorname{tr}F^{ij},$ (4.35)
[…]
$>0$
Second, we choose $B_{A}$ large, depending on $A$, so that
Fix $l\in\{1,2,\dots,n-1\}$, it follows from (4.24) that
Combining (4.20) and (4.21) yields
which implies
by taking $B_{A}$ large. It follows from (4.25) that
Then using (4.37) and (4.38), we have
[…]
By the comparison principle,
Since $\Psi^{l}(0)=0$, we get $\Psi^{l}_{n}(0)\leq 0$, which yields
$|v_{nl}(0)|\leq C,\quad\forall 1\leq l\leq n-1,$
$C\geq M^{2}\left|\tilde{\nu}^{\prime}D^{2}_{y}u(y_{0})\tilde{\eta}^{l}\right|$ $=M^{2}\left|\big(T^{-1}\tilde{\nu}\big)^{\prime}D^{2}_{x}u(x_{0})T^{-1}\tilde{\eta}^{l}\right|$
$=M^{2}\left|u_{\nu\eta^{l}}(x_{0})\right|\frac{1}{|T\nu|}\frac{1}{M_{l}}$
Theorem 4.6.
$\sup_{\partial\Omega}|u_{\nu\nu}|\leq C,$
Proof.
At the origin, we have
$u_{ni}(0)\leq C\sqrt{b_{i}}\quad\text{for}\ i=1,2,\dots,n-2.$ (4.39)
Then it follows from Lemma 4.2, Lemma 4.3, and (4.39) that
Combining Lemmas 3.1, 3.2, 3.3, and Theorem 4.6, we obtain the $C^{2}$ estimate independent of $\inf_{\Omega_{0}}f$, and then Theorem 1.1 follows from an approximation argument.
@@ 5. The weakly interior estimate
$\hskip 12.80373pt+\frac{1}{(\psi+\beta)^{\alpha+1}}\left(\frac{\alpha-1}{4}|\xi|^{2}\psi_{x_{i}}\psi_{x_{j}}-2\alpha\psi_{\xi}\psi_{\xi x_{i}}\psi_{x_{j}}\right)$
$\hskip 12.80373pt+\frac{1}{(\psi+\beta)^{\alpha+1}}\left(-2\alpha\psi_{\xi}\psi_{x_{i}}\psi_{\xi x_{j}}-\alpha\psi_{\xi}^{2}\psi_{x_{i}x_{j}}\right)$
@@
$\hskip 12.80373pt+\frac{G^{ij}}{(\psi+\beta)^{\alpha+1}}\left(\frac{\alpha-1}{4}|\xi|^{2}\psi_{x_{i}}\psi_{x_{j}}-(1+\alpha)\psi_{\xi}\psi_{x_{i}}\xi_{j}\right)$
$\hskip 12.80373pt+\frac{G^{ij}}{(\psi+\beta)^{\alpha+1}}\left(3\alpha\psi_{\xi}^{2}\psi_{x_{i}x_{j}}+\frac{1}{2}\alpha\psi_{\xi}^{2}\delta_{ij}\right)$
@@
$\leq\frac{1}{(\psi+\beta)^{\alpha}}\left(-\frac{1-\alpha}{4\alpha}+C_{1}\|\psi\|_{C^{3}(\overline{\Omega})}^{2}\right)|\xi|^{2}$ (5.10)
$\hskip 12.80373pt+\frac{G^{ij}}{(\psi+\beta)^{\alpha}}\left[-\frac{1}{4\alpha}\xi_{i}\xi_{j}-\frac{1+\alpha}{\psi+\beta}\psi_{\xi}\psi_{x_{i}}\xi_{j}-\frac{\alpha^{2}+3\alpha}{(\psi+\beta)^{2}}\psi_{\xi}^{2}\psi_{x_{i}}\psi_{x_{j}}\right],$
@@
$\hskip 12.80373pt\mathscr{L}\big((\psi+\beta)\eta^{2}\big)$
$=G^{ij}\left(\partial_{x_{i}x_{j}}+2q_{j}\partial_{\eta x_{i}}+q_{i}q_{j}\partial_{\eta\eta}\right)\left[(\psi+\beta)\eta^{2}\right]$
@@
$\hskip 12.80373pt\mathscr{L}(w)$
$=G^{ij}\left(\partial_{x_{i}x_{j}}+2r\partial_{x_{i}\xi_{j}}+2q_{j}\partial_{x_{i}\eta}+r^{2}\partial_{\xi_{i}\xi_{j}}+2rq_{j}\partial_{\xi_{i}\eta}+q_{i}q_{j}\partial_{\eta\eta}-2q_{j}\partial_{\xi_{i}}\right)[w]$
@@
$\geq-K_{1}\sigma_{1}^{\frac{-1}{k-1}}\left(6|\xi|^{2}+\eta^{2}+r^{2}\right)-K_{1}\left(4\eta^{2}+r^{2}\right)-4r^{2}$
$\hskip 12.80373pt-\left(\|u\|_{C^{1}(\overline{\Omega})}^{2}+2\|u\|_{C^{0}(\overline{\Omega})}\right)|q|^{2}$
$\geq-K_{1}\sigma_{1}^{\frac{-1}{k-1}}\left(6|\xi|^{2}+\eta^{2}+\frac{C_{\psi}}{\beta^{2}}|\xi|^{2}\right)-K_{1}\left(4\eta^{2}+\frac{C_{\psi}}{\beta^{2}}|\xi|^{2}\right)$
$\hskip 12.80373pt-4\frac{C_{\psi}}{\beta^{2}}|\xi|^{2}-\left(\|u\|_{C^{1}(\overline{\Omega})}^{2}+2\|u\|_{C^{0}(\overline{\Omega})}\right)\frac{C_{\psi}}{\beta^{4}}|\xi|^{2}$
@@
$\sigma_{1}\big[\lambda\big(D^{2}u\big)\big]\geq\lambda_{\operatorname{max}}\big(D^{2}u\big)\geq u_{\xi_{*}\xi_{*}}\geq\frac{1}{\beta^{3}}-3\|u\|_{C^{1}(\overline{\Omega})}>1,$ (5.19)
@@
$=\left(\frac{1}{4\alpha}(\psi+\beta)^{1-\alpha}+\frac{\psi_{\xi}^{2}}{(\psi+\beta)^{\alpha}}\right)\left(\beta^{\alpha}\sup_{\partial\Omega}u_{\nu\nu}+\frac{4K+2C}{\beta^{8}}\right)$
@@
using the results in Section 3, we complete the proof of Theorem 1.2. ∎